OpenStudy (anonymous):

How would you go about differentiating 1/sqrt[3]{x} (x^3+tan(x))^4. I said u=1/sqrt[3]{x}(x^3+tan(x)) And y=u^4. Can someone tell me why I'm wrong? (the sqrt[3]{x} is suppose to signify cubed root of)

OpenStudy (homeworksucks):

I have no idea how do this, I'm just going to put it in equation for for you

OpenStudy (homeworksucks):

\[\frac{ 1 }{ \sqrt{3} x}(x ^{3}+\tan(x))^{4}\]

OpenStudy (homeworksucks):

Is that right?

OpenStudy (anonymous):

it's a cubed root and the (x^3+tan(x))^4 is within the cubed root

OpenStudy (homeworksucks):

Got it, \[1/\sqrt[3]{x(x^3+\tan(x)^{4}}\]?

OpenStudy (anonymous):

Sorry! But it's without the x out in front of the (X^3...)^4

OpenStudy (homeworksucks):

Oh, so just the x cubed plus tan(x) to the 4th

OpenStudy (homeworksucks):

\[1/\sqrt[3]{(x ^{3}+\tan(x))^{4}}\]

OpenStudy (anonymous):

Yes! My mistake. That is correct.

OpenStudy (homeworksucks):

Cool! I'll let a person who actually knows a shred of calculus handle it from here

OpenStudy (anonymous):

okay, thanks!

zepdrix (zepdrix):

\[\Large \cfrac{1}{\sqrt[3]{(x^3+\tan x)^4}}\]We need derivative? :o So you were making a substitution to simplify it down or something?

OpenStudy (perl):